OpenStudy (anonymous):

Identify the solution(s) of the system of equations, if any. x + 2y = 5 3x - 4y = 0

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

oh wait x = (4/3)*y sub into eq 1 (10/3)*y= 5 so y = 3/2 and x = 2

OpenStudy (jdoe0001):

if you solve for "y" on each, what do you get?

OpenStudy (anonymous):

If we solve using elimination, we get:  5x=10 x = 2 You can plug x=2 into whatever to find y.